OpenStudy (landon34):

Multiply: (-12/25)(-10/16)

OpenStudy (landon34):

I need help.

OpenStudy (anonymous):

both numbers are negative will the answer be positive or negative?

OpenStudy (landon34):

positive.

OpenStudy (anonymous):

ok right, so we can ignore the signs because we know that we have \[\frac{12}{25}\times \frac{10}{16}\] before multiplying do you see any common factors top and bottom?

OpenStudy (anonymous):

for example, what number goes in to both 10 and 25 evenly?

OpenStudy (landon34):

5

OpenStudy (anonymous):

right so we can cancel a 5 top and bottom as a start \[\frac{12}{\cancel{25}^5}\times \frac{\cancel{10}^2}{16}=\frac{12\times 2}{5\times 16}\]

OpenStudy (anonymous):

how about 12 and 16? what is the largest number that divides both of those evenly?

OpenStudy (landon34):

4

OpenStudy (anonymous):

ok so we can cancel some more \[\frac{\cancel{12}^3\times 2}{5\times\cancel{ 16}^4}=\frac{3\times 2}{5\times 4}\]

OpenStudy (anonymous):

and now we can still cancel even more

OpenStudy (anonymous):

because 2 goes in to both 2 and 4 evenly

OpenStudy (landon34):

yes.

OpenStudy (anonymous):

cancel those, and you are left only with \[\frac{3}{5\times 2}\]